Submarine X23 has parametric equation of 2+t= x(t) & 4-3t= y(t) and Submarine Y18 has parametric equation of 11-t= x(t) & 3+at= y(t). how can i find the time they intersect if Submarine X23 left at 2.17 PM ?
Answer by Alan3354(69443)

11-t= x(t) & 3+at= y(t).
You'll have to define a